OpenStudy (anonymous):

can someone explain to me how b sqrt t+(a+bt)/2 sqrt t equates to a+3bt/2 sqrt t

OpenStudy (vijay):

hey, can u write question again with the brackets ?

OpenStudy (anonymous):

can some explain to me how [b sqrt t+(a+bt)]/[2 sqrt t] equate to or simplifies to [a+3bt]/[2 sqrt t]

OpenStudy (vijay):

your question is wrong, plz check again ?

OpenStudy (vijay):

is that your question ? \[[b \sqrt{t}+(a+bt)/2\sqrt{t}] = [a+3bt]/2\sqrt{t}\]

OpenStudy (anonymous):

yes the first part is the answer but the second part is what it is reduced to which is supposed to be the final answer but i dont understnd how the second part is the final answer

OpenStudy (vijay):

actually the right hand side is wrong, that is not equal to left hand side.

OpenStudy (anonymous):

ok